byeeee:

Simplify square root of 2 multiplied by the cube root of 2. 2 to the power of 1 over 6 2 to the power of 2 over 3 2 to the power of 5 over 6 2 to the power of 7 over 6

563blackghost:

So your equation is: \(\large\bf{\sqrt{2} \times \sqrt[3]{2}}\) You would need to change these into rational exponents. The format is: \(\large\bf{\sqrt[n]{x^{m}}}\) So you re-write the first one as: \(\large\bf{2^{\frac{1}{2}}}\) and the second one as: \(\large\bf{2^{\frac{1}{3}}}\) So we now have: \(\large\bf{2^{\frac{1}{2}} \times 2^{\frac{1}{3}}}\) What do we do next?

BenLindquist:

multiply

byeeee:

you have to multiply

byeeee:

it equals 4

byeeee:

??

byeeee:

wait

byeeee:

it equals 7 square root of 2 over 3

563blackghost:

Well yes we need to multiply, `BUT` the exponents need to have the same denominator. How would you do this?

byeeee:

ohhhh wait

byeeee:

i remember learning it but i forgot how you do it omg

563blackghost:

XD okie well to have the same denominator you need to think of a term that has a factor of 2 and 3. Which comes to mind?

byeeee:

1

563blackghost:

Not quiet. How about 6? 2 x 3 = 6 so it definitely works.

byeeee:

ohhhhhhhh

byeeee:

im so stupid

563blackghost:

So we multiply the fractions to get a denominator of 6. \(\large\bf{\frac{1}{2} \times 3 = \frac{3}{6}}\) \(\large\bf{\frac{1}{3} \times 2 = \frac{2}{6}}\) So now you have... \(\large\bf{2^{\frac{3}{6}} \times 2^{\frac{2}{6}}}\)

563blackghost:

What is your answer?

byeeee:

the answer i got isn't an answer choice :/

563blackghost:

You just need to add the exponents together, the 2 stays as 2. \(\Large\bf{\color{red}{2^{\frac{3}{6} + \frac{2}{6}}}}\)

byeeee:

2 and 5 over 6

563blackghost:

yup

byeeee:

yayyyy thank you so much! are you good with graphing stuff

563blackghost:

np kind of >.> but i have to go sadly ;-;

byeeee:

oh okay well thank for your help!